# Optimizing Multi-Hop Document Retrieval Through Intermediate Representations

### v1

@@ Abstract
Jiaen Lin Independent Researcher lje23@mails.tsinghua.edu.cn                        Jingyu Liu Renmin University of China liujy1016@ruc.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have made significant advancements in natural language processing, demonstrating exceptional performance across a variety of tasks Muennighoff et al. (2023); Hendrycks et al. (2021), including complex multi-hop tasks. However, despite their success, LLMs often generate factually incorrect answers as they rely solely on their internal parametric memory, which struggles with tasks requiring external knowledge Lewis et al. (2020). To address this limitation, Retrieval-Augmented Generation (RAG) frameworks Gao et al. (2023) have been developed, utilizing semantic similarity metrics to retrieve relevant external documents, which are subsequently integrated into the generation context. Nevertheless, RAG still faces challenges, particularly in handling complex queries that require multi-hop document retrieval.
@@ 3.3 Layer-wise Dynamics of Transformation Divergence
These three operations reflect fundamental information processing mechanisms within transformer architectures. During the initial high ${\mathrm{TD}({\bm{W}}_{v})}$ phase, LLMs extract information from disorganized data. The subsequent low ${\mathrm{TD}({\bm{W}}_{v})}$ phase signifies the processing of this information. Finally, the resurgent high ${\mathrm{TD}({\bm{W}}_{v})}$ phase aligns the representation with the token prediction requirements which is also proved in Fan et al. (2024) and Haviv et al. (2023). Based on these operations, we can categorize the representations during inference into four types:
@@ 4.2 Implementation Details
The overall contrastive InfoNCE loss for the set of intermediate representations $R$ is computed as the average loss across all representations $r\in R$. It is formally defined as:
@@
where $d_{i}^{+}$ denotes the positive document for representation $r_{i}$ in $D$, and ^$\hat{\mathcal{L}}(r_{i},d^{+}_{i})$ is the contrastive InfoNCE loss defined in Equation 4.
@@ 5.1 Setup
Baselines. Our baseline models encompass: (1) straightforward methods, specifically No Retrieval, Vanilla RAG, and SelfRAG Asai et al. (2024); (2) advanced multi-step retrieval strategies, namely SelfAsk Press et al. (2023a) and IterRetGen Shao et al. (2023a). All models are executed on the LLama3.1-8b platform.

### v2

@@ Abstract
Jiaen Lin1, Jingyu Liu3, Yingbo Liu1, 2 1School of Software, BNRist, Tsinghua University, 2Beijing Key Laboratory of Industrial Bigdata System and Application, 3Gaoling School of Artificial Intelligence, Renmin University of China, Correspondence: lje23@mails.tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have made significant advancements in natural language processing, demonstrating exceptional performance across a variety of tasks Muennighoff et al. (2023); Hendrycks et al. (2021), including complex multi-hop tasks. However, despite their success, LLMs often generate factually incorrect answers as they rely solely on their internal parametric memory, which struggles with tasks requiring external knowledge Lewis et al. (2020). To address this limitation, Retrieval-Augmented Generation (RAG) frameworks Gao et al. (2023b); Liu et al. (2024) have been developed, utilizing semantic similarity metrics to retrieve relevant external documents, which are subsequently integrated into the generation context. Nevertheless, RAG still faces challenges, particularly in handling complex queries that require multi-hop document retrieval.
@@ 3.3 Layer-wise Dynamics of Transformation Divergence
These three operations reflect fundamental information processing mechanisms within transformer architectures. During the initial low ${\mathrm{TD}({\bm{W}}_{v})}$ phase, LLMs extract information from disorganized data. The subsequent high ${\mathrm{TD}({\bm{W}}_{v})}$ phase signifies the processing of this information. Finally, the resurgent low ${\mathrm{TD}({\bm{W}}_{v})}$ phase aligns the representation with the token prediction requirements which is also proved in Fan et al. (2024) and Haviv et al. (2023). Based on these operations, we can categorize the representations during inference into four types:
@@ 4.2 Implementation Details
The overall InfoNCE loss for the set of intermediate representations $R$ is computed as the average loss across all representations $r\in R$. It is formally defined as:
@@
where $d_{i}^{+}$ denotes the positive document for representation $r_{i}$ in $D$, and ^$\hat{\mathcal{L}}(r_{i},d^{+}_{i})$ is the contrastive InfoNCE loss defined in Equation 4. We train L-RAG separately for each dataset to achieve better performance.
For reasoning, the layer with the minimum TD is a natural starting point for selection. L-RAG would refine the choice by exploring neighboring layers to identify the most suitable one. Specifically, we define the set of candidate layers (CL) as:
$CL(k,n,l)=l-nk,\dots,l,\dots,l+nk,$ (6)
where $n$ represents the number of candidate layers (CL) considered on each side, and $k$ denotes the step size. We then train retrievers using each layer in candidate layers and evaluate their performance. The layer yielding the best results is selected as the optimal one.
@@ 5.1 Setup
Baselines. Our baseline models encompass: (1) straightforward methods, specifically No Retrieval, Vanilla RAG, and HyDE Gao et al. (2023a); (2) advanced multi-step retrieval strategies, namely SelfAsk Press et al. (2023a), IterRetGen Shao et al. (2023a) and IRCoT Trivedi et al. (2023). All models are executed on the LLama3.1-8b platform.
@@ 6 Conclusion
Acknowledgments
This work is primarily supported by the Key Research and Development Program of China from MOST(2024YFB3311901). We would like to express our gratitude to the anonymous ARR February reviewers (Reviewer 27qA, xzJs, VoyJ) and Area Chair TVjr for their valuable feedback and suggestions that helped improve this paper.